The complex numbers z1, z2, z3 satisfying z1-z3/z2-z3 = (1 - v3i)/2 are vertices of triangle which is : a) of area zero b) right angled isosceles c) equilaterald) obtuse angled isosceles

Grade:11

1 Answers

moumi roy
7 years ago
ans(c)
i(pi)/3z3 = (1 - v3i)/2=cos60+isin60=ez1-z3/z2-
z1-z3=ei(pi)/3{z2-z3)
hence z1,z2.z3 r pts of equilatral tiangle